Write the standard form of the equation of each line given the slope and y-intercept:
Slope=9, y intercept=4
The answer is 9x-y=-4 but how?

Answer by rfer(16322)

slope intcept form
y=mx+b
y=9x+4
change to standard form
-9x+y=4
multiply through by -1
9x-y=-4